# Stochastic Subspace Descent Accelerated via Bi-fidelity Line Search

### v1

@@ 4.2.1 Dual Form of Kernel Ridge Regression
where $\bm{K}$ is the kernel matrix with $[\bm{K}]_{i,j}=\kappa(\bm{x}_{i},\bm{x}_{j})$, $[\bm{y}]_{i}=y_{i}$, and $\lambda$ is a positive scalar regularization parameter. The solution of Equation (4.3) can be explicitly represented as
@@
We pretend the problem in a black-box format, where access to the HF function is only available through an API, thus hiding $\bm{y}$ (and/or $\bm{K}$) in Equation (4.3) and making derivative information inaccessible. In this case, we assume the values of $\bm{y}$ are unavailable for privacy reasons. For the regression data, we select the first $D=1,\!000$ samples from the California housing dataset provided in the scikit-learn library (Pedregosa et al., 2011). A subset $\mathcal{S}$ with size $l=10$ is randomly selected for the Nyströem approximation. We use a Gaussian (RBF) kernel with lengthscale $1.0$ to generate the corresponding kernel matrix $\bm{K}$. Figure 4 shows the decay of eigenvalues for $\bm{K}$, with a rapid drop, especially within the first 100 eigenvalues, due to the Gaussian kernel’s properties. This fast decay motivates our focus on cases where the values of $\ell$ are below 100. The starting point $\bm{x}_{0}$ is set at the origin $\bm{0}$.
Figure 5: Similar to Figure 3, we compare the optimizer performances with varying parameters $\ell=100$ and $c\in\{0.9,0.95,0.99\}$. The corresponding results are presented with their titles indicating the specific choices. The shadow regions are the area between the best and the worst behavior over 10 trials.
The results of kernel ridge regression are shown in Figure 5, with values of $c$ chosen from 0.9 and 0.99, and the value of $\ell$ is fixed as 100. According to these results, BF-SSD shows advantages over other methods except HF-SSD in Figure 5. When the backtracking factor $c$ decreases, the step sizes determined by the backtracking method become more conservative, leading to suboptimal results, especially for BF-SSD. We also implement different combinations of $c$ and $\ell$ and collect the SSD performances in Table 3. The results suggest that BF-SSD outperforms other methods for most cases, and larger values of $\ell$ and $c$ improve the performance of BF-SSD.
@@
0.9 10 $3497.69\pm 3.56$ $8.98\pm 0.43$ $23.65\pm 0.83$ $\mathbf{8.77\pm 0.72}$
50 $1016.43\pm 9.26$ $\mathbf{3.96\pm 0.17}$ $22.19\pm 0.66$ $6.49\pm 0.51$
100 $268.71\pm 3.32$ $\mathbf{5.70\pm 0.26}$ $21.81\pm 0.71$ $6.01\pm 0.36$
0.95 10 $3499.15\pm 6.41$ $19.17\pm 0.99$ $23.39\pm 0.72$ $\mathbf{3.21\pm 0.40}$
50 $1009.36\pm 9.14$ $5.52\pm 0.13$ $21.46\pm 0.55$ $\mathbf{2.42\pm 0.38}$
100 $269.68\pm 2.79$ $6.07\pm 0.23$ $21.32\pm 1.09$ $\mathbf{2.06\pm 0.27}$
0.99 10 $3499.32\pm 7.90$ $30.18\pm 1.08$ $23.04\pm 0.91$ $\mathbf{1.59\pm 0.74}$
50 $1010.93\pm 6.68$ $6.94\pm 0.21$ $21.94\pm 0.50$ $\mathbf{0.88\pm 0.49}$
100 $270.82\pm 4.04$ $6.17\pm 0.19$ $21.53\pm 0.75$ $\mathbf{0.75\pm 0.38}$
@@ 4.2.2 Black-box Adversarial Attack on MNIST
In this study, we utilized two convolutional neural network (CNN) architectures to model the HF and LF representations for classification tasks on the MNIST dataset with $60,\!000$ training data and $10,\!000$ testing data. The HF model was a deeper CNN consisting of two convolutional layers, the first with 32 filters and the second with 64 filters, both using $5\times 5$ kernels, followed by ReLU activations and $2\times 2$ max-pooling. The flattened output from the convolutional layers ($7\times 7\times 64$) was connected to a fully connected layer with 1024 neurons, followed by a 10-class output layer. In contrast, the LF model employed a simplified architecture with a single convolutional layer containing 2 filters and a $3\times 3$ kernel, followed by ReLU activation and $2\times 2$ max-pooling. The output ($13\times 13\times 2$) was flattened and passed through a fully connected layer with 16 neurons, leading to a 10-class output layer with log-softmax activation. The HF model was designed to provide high-capacity representations, while the LF model served as a lightweight alternative for computational efficiency. The LF model was trained using knowledge distillation (Hinton et al., 2015), leveraging only 1000 training samples and 1000 evaluations of the HF function. Knowledge distillation is a technique where a smaller, simpler model (the student) learns to replicate the outputs of a larger, more complex model (the teacher), effectively transferring knowledge while reducing computational costs. The classification accuracy for the HF and LF CNNs are $99.02\%$ and $82.21\%$, respectively. There are $27,\!562$ parameters for the LF CNN and $3,\!274,\!634$ parameters for the HF CNN. We estimate the ratio between HF and LF computational costs as $3274634/27562\approx 118.8$. The images in MNIST dataset are $28\times 28$ with a single channel, hence the dimension is $D=784$. The starting points are initialized as the origin point for all experiments.
@@
Figure 7: Adversarial examples for MNIST sample #8 (top two rows) and MNIST sample #18 (bottom two rows) using different methods at $N=2,\!000,5,\!000,$ and $7,\!000$.
Figure 6 illustrates the convergence of various zeroth-order methods on two test images. For the SSD methods (including the line search version), the parameters are set to $\ell=50$ and $\alpha_{\textup{max}}=2.0$. Since BF-SSD uses 1,000 HF evaluations for knowledge distillation training, it begins at $N=1,\!000$. The convergence results demonstrate that HF-SSD outperforms other methods in this task. Additionally, HF-SSD, BF-SSD, and SPSA exhibit clear advantages over other methods, underscoring the importance of tuning suitable step sizes for the optimization process.
In Figure 7, we present the adversarially attacked test images generated by different optimization approaches for $N=2,\!000$, $5,\!000$, and $7,\!000$. For the first test image (a-f), only HF-SSD, BF-SSD, and SPSA successfully flip the output of the HF model under limited HF evaluations ($N\leq 7,\!000$). Similarly, for the second test image (g-l), HF-SSD, BF-SSD, SPSA, and VR-SSD succeed in flipping the HF model output. However, in both cases, we observe that HF-SSD (Figure 7 and Figure 7) and BF-SSD (Figure 7 and Figure 7) tend to blur the images more than SPSA (Figure 7 and Figure 7). This behavior may result from differences in sampling strategies, such as Haar measure sampling versus Hadamard sampling. From an adversarial attack perspective, a successful attack should flip the model’s output without excessively blurring the image. In this regard, SPSA outperforms HF-SSD and BF-SSD, although its loss function remains higher than that of the other methods. The reason for this observation requires further investigation but is out of the scope of this work.
@@ 4.2.3 Soft Prompting Black-box Language Model
For the pre-trained classifier and tokenizer, we employed the BERT model, a state-of-the-art transformer-based architecture trained on large corpora. Specifically, we focused on a simplified version of BERT, named DistilBERT. Sentiment analysis on the aclImdb dataset was used as a soft prompting task. This dataset comprises movie reviews categorized into positive and negative sentiments, forming a binary classification problem. A $D=768$ dimension soft prompt $\bm{x}$ is considered as the input. The transformer’s parameters were kept frozen to focus optimization on the soft prompt, reducing the degrees of freedom and computational overhead. For computational convenience, we focus on small-scale in this example. The HF model, as described in Equation (4.7), was evaluated using 10 samples from aclImdb to approximate the expectation, while the LF model leveraged only 2 samples that were randomly selected from them. Consequently, the evaluation cost ratio between HF and LF was 5:1.
We set the initial starting point at the origin. We let $\ell=50$, $c=0.99$, and for the methods without line search we chose a fixed step size of $1\times 10^{-2}$. The $y$-axis in the following figure represents the average cross entropy loss. Figure 8 illustrates the performances of different competing methods with the BF-SSD demonstrating notable advantages.
@@ 5 Conclusion
In this work, we proposed and analyzed a bi-fidelity line search scheme designed to accelerate the convergence of zeroth-order optimization algorithms. By constructing a surrogate model using both computationally expensive high-fidelity (HF) and inexpensive low-fidelity (LF) objective function evaluations, our approach enables an efficient backtracking line search on the surrogate to determine suitable step sizes. This significantly reduces the reliance on costly HF evaluations, a common bottleneck in zeroth-order methods. We established theoretical convergence guarantees for this scheme under standard assumptions and subsequently integrated it into the stochastic subspace descent framework, yielding the Bi-Fidelity Stochastic Subspace Descent (BF-SSD) algorithm.
@@
Acknowledgment
This work was supported by the AFOSR awards FA9550-20-1-0138, with Dr. Fariba Fahroo as the program manager, and by the US Department of Energy’s Wind Energy Technologies Office.
@@ B.1 Assuming Strong-convexity
Define the filtration $\mathcal{F}_{k}\coloneqq\sigma(\bm{P}_{1},\dots,\bm{P}_{k-1})$ and $\mathcal{F}_{1}=\{\emptyset,\Omega\}$. By Lipschitz continuity, we have
@@
[…]
@@
According to Equation (B.3), the exact line search method can find the optimal $\alpha_{k}$ such that the quadratic term in Equation (B.2) yields $-\alpha_{k}+D\alpha_{k}^{2}L/2\ell\leq-\ell/(2dL)$ for any $\bm{P}_{k}$, thereby
@@
(Showing the existence of a feasible set such that the Armijo condition is satisfied.) Following Equation (B.4), the backtracking method selects the maximal possible step size value that satisfies the Armijo condition with specified parameter $\beta\leq\ell/2d$ and shrinking parameter $c<1$. When $0\leq\alpha_{k}\leq\ell/DL$, $-\alpha_{k}+D\alpha_{k}^{2}L/2\ell\leq-\alpha_{k}/2$ holds with Haar measure probability one, which implies the following Armijo stopping condition
@@
Since $00<c<$ 0, $0<\gamma\leq L$, $0<\ell\leq D$, and $0.50<\beta<$ 0.5, the term 0<($10<\left(1-\min\{2\gamma\beta,2\ell c\gamma\beta/DL\}\right)<$ 1, the convergences $f(\bm{x}_{k})\overset{a.s.}{\to}f^{*}$ and $f(\bm{x}_{k})\overset{L^{1}}{\to}f^{*}$ are guaranteed. ∎
@@ Appendix C Worst Function in the World: Additional Data
Table 4: Performance values for different optimization methods across various $c$ and $\ell$ combinations at $N=5,\!000$. The minimum value in each row is highlighted in bold.

### v2

@@ 4.2.1 Dual Form of Kernel Ridge Regression
where $\bm{K}$ is the kernel matrix with $\bm{K}]_{i,j}=\kappa(\bm{x}_{i},\bm{x}_{j})$, $\bm{y}]_{i}=y_{i}$, and $\lambda$ is a positive scalar regularization parameter. The solution of Equation (4.3) can be explicitly represented as
@@
We pretend the problem in a black-box format, where access to the HF function is only available through an API, thus hiding $\bm{y}$ (and/or $\bm{K}$) in Equation (4.3) and making derivative information inaccessible. In this case, we assume the values of $\bm{y}$ are unavailable for privacy reasons. For the regression data, we select the first D=1, [^0] D=1,\!000 samples from the California housing dataset provided in the scikit-learn library (Pedregosa et al., 2011). A subset $\mathcal{S}$ with size $l=10$ is randomly selected for the Nyströem approximation. We use a Gaussian (RBF) kernel with lengthscale $1.0$ to generate the corresponding kernel matrix $\bm{K}$. Figure 4 shows the decay of eigenvalues for $\bm{K}$, with a rapid drop, especially within the first 100 eigenvalues, due to the Gaussian kernel’s properties. This fast decay motivates our focus on cases where the values of $\ell$ are below 100. The starting point $\bm{x}_{0}$ is set at the origin $\bm{0}$.
Figure 5: Similar to Figure 3, we compare the optimizer performances with varying parameters $\ell=100$ and $c\in\{0.9,0.99\}$. The corresponding results are presented with their titles indicating the specific choices. The shadow regions are the area between the best and the worst behavior over 10 trials.
The results of kernel ridge regression are shown in Figure 5, with values of $c$ chosen from 0.9 and 0.99, and the value of $\ell$ is fixed as 100. According to these results, BF-SSD shows advantages over other methods except NAG and HF-SSD in Figure 5. When the backtracking factor $c$ decreases, the step sizes determined by the backtracking method become more conservative, leading to suboptimal results, especially for BF-SSD. We also implement different combinations of $c$ and $\ell$ and collect the SSD performances in Table 3. The results suggest that BF-SSD outperforms other SSD methods for most cases, and larger values of $\ell$ and $c$ improve the performance of BF-SSD.
$c$ $\ell$ FS-SSD HF-SSD VR-SSD BF-SSD
0.9 10 $3500.91\pm 5.74$ $9.49\pm 0.21$ $23.31\pm 0.83$ $\mathbf{8.18\pm 0.60}$
50 $1015.82\pm 5.90$ $\mathbf{4.01\pm 0.13}$ $20.87\pm 0.55$ $6.13\pm 0.22$
100 $270.95\pm 7.07$ $\mathbf{5.49\pm 0.18}$ $21.56\pm 0.69$ $6.22\pm 0.25$
0.95 10 $3493.88\pm 11.27$ $18.39\pm 0.28$ $23.20\pm 0.30$ $\mathbf{2.94\pm 0.49}$
@@
[…]
0.99 10 $3503.99\pm 3.77$ $29.55\pm 0.05$ $23.35\pm 0.38$ $\mathbf{1.40\pm 0.29}$
50 $1010.92\pm 7.17$ $6.95\pm 0.19$ $20.90\pm 0.31$ $\mathbf{0.75\pm 0.12}$
100 $273.64\pm 3.41$ $6.26\pm 0.08$ $21.38\pm 0.46$ $\mathbf{0.79\pm 0.24}$
@@ 4.2.2 Black-box Adversarial Attack on MNIST
In this study, we utilized two convolutional neural network (CNN) architectures to model the HF and LF representations for classification tasks on the MNIST dataset with 60,00060,\!000 training data and 10,00010,\!000 testing data. The HF model was a deeper CNN consisting of two convolutional layers, the first with 32 filters and the second with 64 filters, both using $5\times 5$ kernels, followed by ReLU activations and $2\times 2$ max-pooling. The flattened output from the convolutional layers ($7\times 7\times 64$) was connected to a fully connected layer with 1024 neurons, followed by a 10-class output layer. In contrast, the LF model employed a simplified architecture with a single convolutional layer containing 2 filters and a $3\times 3$ kernel, followed by ReLU activation and $2\times 2$ max-pooling. The output ($13\times 13\times 2$) was flattened and passed through a fully connected layer with 16 neurons, leading to a 10-class output layer with log-softmax activation. The HF model was designed to provide high-capacity representations, while the LF model served as a lightweight alternative for computational efficiency. The LF model was trained using knowledge distillation (Hinton et al., 2014), leveraging only 1000 training samples and 1000 evaluations of the HF function. Knowledge distillation is a technique where a smaller, simpler model (the student) learns to replicate the outputs of a larger, more complex model (the teacher), effectively transferring knowledge while reducing computational costs. The classification accuracy for the HF and LF CNNs are $99.02\%$ and $82.21\%$, respectively. There are 27,56227,\!562 parameters for the LF CNN and 3,274,6343,\!274,\!634 parameters for the HF CNN. We estimate the ratio between HF and LF computational costs as $3274634/27562\approx 118.8$. The images in MNIST dataset are $28\times 28$ with a single channel, hence the dimension is $D=784$. The starting points are initialized as the origin point for all experiments.
@@
Figure 7: Adversarial examples for MNIST sample #8 (top two rows) and MNIST sample #18 (bottom two rows) using different methods at N=2,000,5,000,N=2,\!000,5,\!000, and 7,0007,\!000.
Figure 6 illustrates the convergence of various zeroth-order methods on two test images. For the SSD methods (including the line search version), the parameters are set to $\ell=50$ and $\alpha_{\textup{max}}=2.0$. Since BF-SSD uses 1,000 HF evaluations for knowledge distillation training, it begins at N=1, [^0] N=1,\!000. The convergence results demonstrate that HF-SSD outperforms other methods in this task. Additionally, HF-SSD, BF-SSD, and SPSA exhibit clear advantages over other methods, underscoring the importance of tuning suitable step sizes for the optimization process.
In Figure 7, we present the adversarially attacked test images generated by different optimization approaches for N=2, [^0] N=2,\!000, 5,0005,\!000, and 7,0007,\!000. For the first test image (a-f), only HF-SSD, BF-SSD, and SPSA successfully flip the output of the HF model under limited HF evaluations ($N\leq 7,\!000$). Similarly, for the second test image (g-l), HF-SSD, BF-SSD, SPSA, and VR-SSD succeed in flipping the HF model output. However, in both cases, we observe that HF-SSD (Figure 7 and Figure 7) and BF-SSD (Figure 7 and Figure 7) tend to blur the images more than SPSA (Figure 7 and Figure 7). This behavior may result from differences in sampling strategies, such as Haar measure sampling versus Hadamard sampling. From an adversarial attack perspective, a successful attack should flip the model’s output without excessively blurring the image. In this regard, SPSA outperforms HF-SSD and BF-SSD, although its loss function remains higher than that of the other methods. The reason for this observation requires further investigation, but it is out of the scope of this work.
@@ 4.2.3 Soft Prompting Black-box Language Model
For the pre-trained classifier and tokenizer, we employed the BERT model, a state-of-the-art transformer-based architecture trained on large corpora. Specifically, we focused on a simplified version of BERT, named DistilBERT. Sentiment analysis on the aclImdb dataset was used as a soft prompting task. This dataset comprises movie reviews categorized into positive and negative sentiments, forming a binary classification problem. A $D=768$ dimension soft prompt $\bm{x}$ is considered as the input. The transformer’s parameters were kept frozen to focus optimization on the soft prompt, reducing the degrees of freedom and computational overhead. For computational convenience, we focus on small scale in this example. The HF model, as described in Equation (4.7), was evaluated using 10 samples from aclImdb to approximate the expectation, while the LF model leveraged only 2 samples that were randomly selected from them. Consequently, the evaluation cost ratio between HF and LF was 5:1.
We set the initial starting point at the origin. We let $\ell=50$, $c=0.99$, and for the methods without line search we chose a fixed step size of $1\times 10^{-2}$. The $y$-axis in the following figure represents the average cross-entropy loss. Figure 8 illustrates the performances of different competing methods with the BF-SSD demonstrating notable advantages.
@@ 5 Conclusion
In this work, we proposed and analyzed a bi-fidelity line search scheme designed to accelerate the convergence of zeroth-order optimization algorithms. By constructing a surrogate model using both computationally expensive high-fidelity (HF) and inexpensive low-fidelity (LF) objective function evaluations, our approach enables an efficient backtracking line search on the surrogate to determine suitable step sizes. This method significantly reduces the reliance on costly HF evaluations, a common bottleneck in zeroth-order methods. We established theoretical convergence guarantees for this scheme under standard assumptions and subsequently integrated it into the stochastic subspace descent framework, yielding the Bi-Fidelity Stochastic Subspace Descent (BF-SSD) algorithm.
@@ B.1 Assuming Strong-convexity
In particular,
$\mathbb{E}[f(\bm{x}_{k+1})]-f^{*}\leq\omega^{k+1}(f(\bm{x}_{0})-f^{*}),$ (B.1)
for $\omega\in(0,1)$ where
$\omega=\begin{cases}1-\frac{\gamma\ell}{DL}&\text{exact line search}\\ 1-\min\left\{2\gamma\beta\alpha_{\textup{max}},\frac{2\ell c\gamma\beta}{DL}\right\}&\text{backtracking}.\end{cases}$
@@
Define the filtration $\mathcal{F}_{k}\coloneqq\sigma(\bm{P}_{1},\dots,\bm{P}_{k-1})$ and $\mathcal{F}_{1}=\{\emptyset,\Omega_{P}\}$, with $\Omega_{P}$ as the sample space. By Lipschitz continuity, we have
@@
    $\text{s.t.}\quad f(\bm{x}_{k}-c^{m}\alpha_{\textup{max}}\bm{P}_{k}\bm{P}_{k}^{T}\nabla f(\bm{x}_{k}))$ $\leq f(\bm{x}^{k})-\beta c^{m}\alpha_{\textup{max}}\lVert\bm{P}_{k}\bm{P}_{k}^{T}\nabla f(\bm{x}^{k})\rVert^{2}.$
@@
According to Equation (B.4), the exact line search method can find the optimal $\alpha_{k}$ such that the quadratic term in Equation (B.3) yields $\alpha_{k}+D\alpha_{k}^{2}L/2\ell\leq-\ell/(2dL)$ for any $\bm{P}_{k}$, thereby
@@
(Showing the existence of a feasible set such that the Armijo condition is satisfied.) Following Equation (B.5), the backtracking method selects the maximal possible step size value that satisfies the Armijo condition with specified parameter $\beta\leq\ell/2d$ and shrinking parameter $c<1$. When $0\leq\alpha_{k}\leq\ell/DL$, $\alpha_{k}+D\alpha_{k}^{2}L/2\ell\leq-\alpha_{k}/2$ holds with Haar measure probability one, which implies the following Armijo stopping condition
@@
Since $0<c<0$, $0<\gamma\leq L$, $0<\ell\leq D$, and $0<\beta<0.5$, the term $0<\left(1-\min\{2\gamma\beta,2\ell c\gamma\beta/DL\}\right)<1$, the convergences $f(\bm{x}_{k})\overset{a.s.}{\to}f^{*}$ and $f(\bm{x}_{k})\overset{L^{1}}{\to}f^{*}$ are guaranteed. ∎
@@ Appendix C Worst Function in the World: Additional Data
Table 4: Performance values for different optimization methods across various $c$ and $\ell$ combinations at N=5, [^0] N=5,\!000. The minimum value in each row is highlighted in bold.